### v1

# Graph Signal Processing for Global Stock Market Volatility Forecasting

Abstract
The interconnectedness of global financial markets has brought increasing attention to modeling volatility spillover effects. Via incorporating Graph Signal Processing techniques, a novel multivariate framework, extending the traditional Heterogeneous Auto-Regressive model, is developed in the spectral domain constructed by the graph Fourier transformation method. Further, a set of convolution filters with learnable weights is employed to more flexibly aggregate the past mid-term and long-term information. Using 24 global stock market indices, the effectiveness of the proposed model is demonstrated through comprehensive empirical evaluations.
Keywords: realized volatility forecasting; multivariate time series forecasting; spatio-temporal analysis; graph signal processing.
1 Introduction
In the complex ecosystem of global financial markets, understanding and forecasting volatility is important for investors, risk managers and policymakers. Unlike the retrospective approach of relying on historical volatility, forecasting volatility provides forward-looking estimates that are more reflective of the present market dynamics and potential upcoming changes. Recent studies have emphasized the volatility spillover phenomenon, which refers to the transmission of volatility shocks from one market to others globally (Kanas, 2000; Forbes and Rigobon, 2002; Poon and Granger, 2003; Diebold and Yilmaz, 2009; Yang and Zhou, 2017; Bollerslev et al., 2018). This phenomenon has gained much attention, especially after the highly volatile 2008 global financial crisis (GFC) and the recent COVID-19. The volatility spillover effect highlights the intricacies of the international financial ecosystem, which necessitates a comprehensive framework to consider the relationships between stock markets when conducting volatility forecasting.
This research aims to advance the field of volatility forecasting by developing a multivariate model that integrates the volatility spillover effect across multiple stock markets, via employing the Graph Signal Processing (GSP) technique (Chen et al., 2015; Wu et al., 2021). GSP extends traditional signal processing to data defined on graphs. A graph can be defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes representing the entities in the graph and $\mathcal{E}$ is the set of edges representing the relationships or links between nodes. GSP provides a powerful framework for analyzing and manipulating complex networks and relational data. This study adapts the GSP tools such as the Graph Fourier Transformation (GFT) and spectral analysis techniques to volatility data. The proposed framework and the empirical findings are expected to help financial practitioners better understand the intricate behaviors of global market volatility.
The organization of this paper is as follows. Section 2 provides a review of pertinent literature. Section 3 discusses the motivation for this study. Section 4 details the methods and practical application of the proposed model. Section 5 presents the conducted experiments and their results and the corresponding codes are available at https://github.com/MikeZChi/GSPHAR.git. Section 6 concludes the discussion.
2 Literature Review
This review section aims to provide a clear understanding of how volatility is measured and to examine different univariate and multivariate models used in forecasting volatility.
2.1 Volatility Measurement
Although volatility has been regarded as an unobservable and latent factor behind the fluctuations in an asset’s returns over a given time period, various methods for measuring volatility have been suggested in the literature (Poon and Granger, 2003). Among different measurement approaches, the high-frequency-data-based Realized Volatility (RV), proposed by Andersen and Bollerslev (1998), is used by this research to represent the volatility of different stock markets. The RV assesses the volatility of an asset at time $t$ by tracking its return over a sample period using a sufficiently high frequency of sampling. This measure is particularly useful for capturing the day-to-day fluctuations in a market index and is extensively used in financial practice. Hence, RV is the forecast target of the proposed model and is reviewed in this section.
With access to high-frequency data, RV represents the daily volatility for day $t$ because it is derived from the sum of squared high-frequency intraday returns, as defined below:
$\text{RV}_{t}=\sum_{m=1}^{M}r_{m,t}^{2},$ (1)
where RVt is the RV data on day $t$, the sampling frequency is $\frac{1}{M}$ and $r_{m,t}$ denotes the $m$-th intraday high-frequency return on day $t$.
2.2 Univariate and Multivariate Volatility Forecasting
Since the proposed model is designed to forecast RV time series data, the discussion focuses on models pertinent to RV forecasting problems. Besides, given that the experiments of this research involve forecasting future RV values for stock market indices worldwide, models reviewed in this section target RV forecasting on stock market indices by default. Notably, the Heterogeneous Auto-Regressive (HAR) model, initially introduced by Corsi (2009) and later refined by Bollerslev et al. (2018), stands out as a frequently employed model in RV forecasting. In addition to the univariate HAR model, this section also explores several multivariate extensions of the HAR model.
The univariate HAR model uses the pooled panel data as the pattern of volatility in three previous time intervals to forecast future RV values. The pooled panel data include the past daily RV observation, the average of the past weekly RV values and the average of the past monthly RV values, which represent the short-term, mid-term and long-term volatility behaviors, respectively. Here, RVt denotes the daily RV data of the chosen stock market index on day $t$ calculated as per Equation (1). The univariate HAR model is formulated as:
$\text{RV}_{t}=\alpha+\beta_{d}\text{RV}_{t-1}+\beta_{w}\text{RV}_{t-5:t-1}+\beta_{m}\text{RV}_{t-22:t-1}+\epsilon_{t},\;\;\epsilon_{t}\sim N(0,\sigma^{2}),$ (2)
where $\text{RV}_{t-n:t-1}=\frac{1}{n}\sum_{j=t-n}^{t-1}\text{RV}_{j}$ is the average RV of the stock market index from time $t-1$ to time $t-n$. Here, coefficients $\beta_{d}$, $\beta_{w}$ and $\beta_{m}$ are scalars and they measure how the past daily, weekly and monthly volatility patterns influence future RVs. The intercept term $\alpha$ is also a scalar. Since the original RV data usually shows leptokurtosis (fat tails) and skewness, the log-transformed and square root versions of the RV data are often utilized to improve the forecasting accuracy of the HAR model. However, the univariate HAR model cannot capture the comovement or correlations of the volatility in different stock markets, which significantly limits its ability to deliver more accurate forecasts.
In order to integrate the volatility spillover effect to generate more accurate forecasts, the Vector HAR (VHAR) and the HAR Kitchen Sink (HAR-KS) models are proposed by Bubák et al. (2011) and Liang et al. (2020), respectively. These models align with the general design of the univariate HAR model. They also use past daily RV data as the short-term volatility pattern, past weekly RV data as the mid-term volatility pattern and past monthly RV data as the long-term volatility pattern to forecast future RV values. The VHAR model incorporates lagged daily, weekly and monthly RV data from all other stock market indices to capture the volatility inter-relationships. In contrast, the HAR-KS model only relies on lagged daily RV observations from all indices. The formulation of the VHAR model and the HAR-KS model are presented in Equation (3) and Equation (4), respectively,
$\begin{split}\left(\text{RV}_{i,t}\right)^{1/2}=&\;\beta_{i,0}+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,d}\left(\text{RV}_{k,t-1}\right)^{1/2}+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,w}\left(\text{RV}_{k,t-5:t-1}\right)^{1/2}\\ &\;+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,m}\left(\text{RV}_{k,t-22:t-1}\right)^{1/2}+\varepsilon_{i,t},\end{split}$ (3)
$\begin{split}\left(\text{RV}_{i,t}\right)^{1/2}=&\;\beta_{i,0}+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,d}\left(\text{RV}_{k,t-1}\right)^{1/2}+\beta_{i,w}\left(\text{RV}_{i,t-5:t-1}\right)^{1/2}\\ &\;+\beta_{i,m}\left(\text{RV}_{i,t-22:t-1}\right)^{1/2}+\varepsilon_{i,t},\end{split}$ (4)
where both $i$ and $k$ are indices of different stock market indices taken into consideration. Here, $i$ is used to denote the target stock market indices and $k$ indexes the included lagged RV variables from all stock market indices to account for the volatility spillover effect. The square root lagged daily, weekly or monthly RV values from other stock market indices, apart from the target index, are treated as additional attributes to capture the volatility spillover effect.
The VHAR and HAR-KS models not only inherit the simple architecture of the univariate HAR model, but are also empirically proven to outperform it in multivariate RV forecasting problems. However, they have some common limitations. The first is scalability. When the number of relevant stock market indices $N$ increases, the number of parameters of the VHAR and the HAR-KS models grows quadratically. Thus, both models have insufficient scalability. In addition, the volatility spillover effect can be much more complex (e.g., nonlinear) than the linear combinations of past RV data, which cannot be captured by the linear VHAR and HAR-KS models.
Without dramatically sophisticating the HAR model, the Graph Neural Network (GNN) HAR model is proposed by Zhang et al. (2023) to generate more accurate multivariate RV forecasts. Here, the model is denoted as the GNNHAR model. The GNNHAR model precomputes the relational network which remains fixed during the training process. The precomputed relational network can be represented by a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ as mentioned in the Introduction section. In graph theory, the structure or connectivity of the relational graph with $N$ nodes can be captured by the adjacency matrix $\mathbf{A}=[\mathbf{A}_{ij}]\in\mathbb{R}^{N\times N}$. For a group of $N$ stock markets, Zhang et al. (2023) used the following way to construct a graph. If stock market $i$ and stock market $j$ are found to be correlated with each other, an edge is constructed to link node $i$ and node $j$, and $\mathbf{A}_{ij}=\mathbf{A}_{ji}=1$. Otherwise, there is no edge between node $i$ and node $j$, and $\mathbf{A}_{ij}=\mathbf{A}_{ji}=0$. Correspondingly, the degree matrix $\mathbf{D}$ can be calculated as: $\mathbf{D}=\text{diag}(d_{1},\ldots,d_{N})$, where $d_{i}=\sum_{j=1}^{N}\mathbf{A}_{ij}$.
GNN models can be divided into two categories: spectral and spatial GNNs. Spectral GNNs rely on the spectral graph theory, whereas spatial GNNs depend on message-passing mechanisms (Wu et al., 2021). In the GNNHAR model, spatial GNNs are applied. Spatial GNNs are specialized neural networks for graph data. It captures the complex interactions between nodes through neighborhood information aggregation and transformation. A neighborhood of a specific node refers to the set of nodes directly connected to the particular node. In each layer, the spatial GNN updates the attributes of a specific entity (node) based on the information of its neighborhood. By stacking multiple layers with linear and nonlinear information transformation functions, each entity in the graph can receive processed information from a larger neighborhood. Suppose $\mathbf{v}_{t}$ contains the square root RV value of each of $N$ stock market indices at day $t$: $\mathbf{v}_{t}=[\left(\text{RV}_{1,t}\right)^{1/2},\left(\text{RV}_{2,t}\right)^{1/2},\ldots,\left(\text{RV}_{N,t}\right)^{1/2}]$ and $\mathbf{v}_{t-n:t-m}$ is the average-pooled data: $\mathbf{v}_{t-n:t-m}=\frac{1}{n-m}\sum_{j=t-n}^{t-m}\mathbf{v}_{t}$. The GNNHAR model adds a $l$-layer spatial GNN component to the HAR model as an additional input variable to capture the volatility spillover effect. The model can be formulated as:
$\begin{split}\mathbf{H}^{(0)}=&\;[\mathbf{v}_{t-1},\mathbf{v}_{t-5:t-2},\mathbf{v}_{t-22:t-6}],\\ \mathbf{H}^{(1)}=&\;\text{ReLU}(\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\mathbf{H}^{(0)}\mathbf{W}^{(1)}),\\ &\ldots,\\ \mathbf{H}^{(l+1)}=&\;\text{ReLU}(\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l+1)}),\\ \mathbf{v}_{t}=&\;\bm{\alpha}+\beta_{d}\mathbf{v}_{t-1}+\beta_{w}\mathbf{v}_{t-5:t-2}+\beta_{m}\mathbf{v}_{t-22:t-6}\\ &\;+\bm{\gamma}\mathbf{H}^{(l+1)}+\bm{\epsilon}_{t}.\end{split}$ (5)
Here, $\mathbf{W}^{(l)}$ is the trainable graph information transformation parameter of layer $l$, representing the linear information transformation, and ReLU()($\cdot$)( ⋅ ) function transforms the graph information nonlinearly. $\bm{\gamma}$ is also a trainable parameter to measure the overall influence of related stock market indices. $\beta_{d}$, $\beta_{w}$ and $\beta_{m}$ are scalars measuring the impact of the lagged panel RV observations similar to the HAR model.
Although all the VHAR, HAR-KS and GNNHAR models follow the same general pattern: adding extra input variables to the HAR model to learn the volatility spillover effect and deliver more accurate multivariate RV forecasts, the GNNHAR model uses spatial GNN to capture the relationships between stock market indices in both linear and nonlinear ways. The inclusion of extra GNN inputs equips the GNNHAR model with additional power to capture the nonlinear volatility spillover effect. In addition, the GNNHAR model applies the same set of weights $\mathbf{W}^{(1)}$, $\mathbf{W}^{(2)}$, …, $\mathbf{W}^{(l+1)}$} to different parts of the volatility network to model the volatility spillover effect. This design ensures the GNNHAR model is scalable. Although this seems to bring the GNNHAR model more power to scale up to larger graphs, trying different numbers of layers $l$ to determine the best scope of the volatility spillover influence can be very time-consuming. In addition, the pattern of the volatility spillover effect can vary across time, which reflects different periods of strengthened or weakened correlations (Diebold and Yilmaz, 2012; Antonakakis et al., 2018; Gong et al., 2021). Consequently, the impact range of the volatility spillover effect can also change over time. Thus, the fixed number of layers ($l$) in the GNNHAR model can limit its adaptability and generalizability.
3 Research Motivation
Even though the univariate HAR model and its multivariate extensions have their own constraints, the general architecture of the HAR model is still preferred in the literature. The HAR model is designed to account for the fact that different market participants (such as high-frequency traders, institutional investors, and long-term investors) operate on different time horizons. The HAR model incorporates pooled RV data from these various time scales, typically daily, weekly and monthly. This model design helps to capture the heterogeneous nature of volatility patterns in financial markets. In addition, empirically it achieves good out-of-sample RV forecasting performance with its simple and intuitive structure, which makes it even more attractive to researchers and practitioners.
Based on the discussions in the previous section, the main challenges for multivariate HAR models are their insufficient expressiveness and limited scalability to larger financial systems. These limitations stem from their architectures where each multivariate HAR model tries to account for the volatility spillover effect by including extra input variables in the model. Although the GNNHAR model attempts to address these issues by applying spatial GNNs, the fixed volatility spillover range reduces effectiveness when modeling complex RV dynamics. In addition, incorporating extra input variables to model volatility spillover effectively learns the residuals between the forecast values generated by the univariate HAR model and the true values. However, treating volatility spillover as merely a compensatory mechanism for the univariate HAR model potentially underestimates its significance. According to Poon and Granger (2003), both the interdependent and the autoregressive behaviors of multi-asset volatility dynamics are salient phenomena. Thus, the volatility spillover effect should be as important as the autoregressive relationship inherent in the HAR models, rather than as an additive correction for neglecting cross-entity relationships. The existing multivariate HAR models do not prioritize these volatility inter-relationships, which could potentially limit their overall forecasting capability. Furthermore, the volatility spillover effect is unlikely to remain constant over time, because market conditions, economic factors and investor behavior continuously change. However, few multivariate RV forecasting models adequately account for this inherent dynamic nature.
This research aims to overcome the above constraints and further improve the multivariate RV forecasting accuracy by proposing another multivariate extension of the HAR model. This model relies on GSP techniques, so the model is referred to as the GSPHAR model. The general idea is to build the HAR model in the spectral domain constructed by the GFT method so that the relational information can be well merged with the RV data. Further, a set of convolution filters with learnable weights is employed to more flexibly aggregate the past mid-term and long-term information. The proposed model can also incorporate extra designs to model the dynamic volatility spillover effect. A Demonstration of the proposed framework is shown in Figure 2 in the following methodology section.
4 Methodology and Implementation
Since the proposed model relies on GSP techniques, this section introduces key concepts from GSP and spectral graph theory, with a focus on the GFT and the magnet framework (Zhang et al., 2021; Lin and Gao, 2023). In addition, this section discusses the graph construction method used to model the RV relationship network among different market indices and a simple yet effective dynamic modeling method.
4.1 Spectral Graph Theory and GSP
As briefly discussed in the Introduction and Literature Review sections, a graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the node set and $\mathcal{E}$ is the edge set of the graph. So far we only consider the undirected and unweighted graphs. However, edges between nodes can have attributes such as direction and weights. A directed edge $e_{m}\in\mathcal{E}$ can be formulated as $e_{m}=(v_{\text{start}},v_{\text{end}})$. It starts from node $v_{\text{start}}\in\mathcal{V}$ and ends at node $v_{\text{end}}\in\mathcal{V}$, which points the information flow direction. In contrast, for an undirected edge, information can travel in both directions without any designated starting or end nodes. An undirected edge can also be regarded as a bidirectional edge because $e_{m}=(v_{i},v_{j})=(v_{j},v_{i})$. According to whether the edges are directed or not, graphs can be divided into undirected graphs and directed graphs, as shown in Figure 1 below. Weights are another edge attribute. Weights are usually positive real numbers and they can be interpreted as the importance of the corresponding edges. Depending on whether the edges have weights or not, graphs can be divided into unweighted graphs and weighted graphs.
(a) An example of undirected graph
(b) An example of directed graph
Figure 1: Undirected and directed graphs
The adjacency matrix $\mathbf{A}$ is used to capture the structure of graphs. If graph $\mathcal{G}$ has $N$ nodes, $\mathbf{A}\in\mathbb{R}^{N\times N}$. For an undirected graph, its adjacency matrix is always symmetric because if $(v_{i},v_{j})\in\mathcal{E}$, $(v_{j},v_{i})\in\mathcal{E}$ as well. Elements in $\mathbf{A}$ can reflect the edge weights. The elements in the adjacency matrix of an undirected and unweighted graph can only take values of 00 or $1$ to reflect connectivity, as shown below:
$\mathbf{A}_{ij}=\begin{cases}1,&\mbox{$(v_{i},v_{j})\in\mathcal{E}$};\\ 0,&\mbox{$(v_{i},v_{j})\notin\mathcal{E}$}.\end{cases}$ (6)
However, for a directed and weighted graph, if $e_{m}=(v_{i},v_{j})\in\mathcal{E}$ and the weights attached to edge $e_{m}$ is $w_{ij}$, $\mathbf{A}_{ij}=w_{ij}$, and in general $\mathbf{A}_{ij}\neq\mathbf{A}_{ji}$. The formulation of the adjacency matrix of the directed and weighted graph is presented below:
$\mathbf{A}_{ij}=\begin{cases}w_{ij},&\mbox{$(v_{i},v_{j})\in\mathcal{E}$};\\ 0,&\mbox{$(v_{i},v_{j})\notin\mathcal{E}$}.\end{cases}$ (7)
No matter whether the graph is directed or weighted, the degree matrix $D$ is always formulated as $\mathbf{D}=\text{diag}(d_{1},\ldots,d_{N})$, where $d_{i}=\sum_{j=1}^{N}\mathbf{A}_{ij}$, for a graph $\mathcal{G}$ with $N$ nodes.
The graph Laplacian matrix $\mathbf{L}$ can be formulated based on the adjacency matrix $\mathbf{A}$ and the degree matrix $\mathbf{D}$:
$\mathbf{L}=\mathbf{D}-\mathbf{A}.$ (8)
The normalized graph Laplacian matrix $\widetilde{\mathbf{L}}$ can be formulated as:
$\widetilde{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}.$ (9)
The degree matrix $\mathbf{D}$ is a diagonal matrix and thus always symmetric, while the adjacency matrix $\mathbf{A}$ is only symmetric when the corresponding graph $\mathcal{G}$ is undirected. If both $\mathbf{D}$ and $\mathbf{A}$ are symmetric and elements in $\mathbf{A}$ (i.e., edge weights) are non-negative, the normalized graph Laplacian matrix $\widetilde{\mathbf{L}}$ is also symmetric, positive semidefinite and diagonalizable. Its eigendecomposition can be formulated as:
$\widetilde{\mathbf{L}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ (10)
where the columns of $\mathbf{U}$ are the eigenvectors of $\widetilde{\mathbf{L}}$, the matrix $\bm{\Lambda}$ is a diagonal matrix containing the corresponding non-negative eigenvalues of $\widetilde{\mathbf{L}}$ and $\mathbf{U}^{\top}$ is the transpose of $\mathbf{U}$. Especially, the eigenvectors in $\mathbf{U}$, which are referred to as the graph Fourier bases, can form a set of orthonormal bases and the eigenvalues are arranged in ascending order from the top left element to the bottom right element in $\bm{\Lambda}$. Large eigenvalues indicate that the corresponding eigenvectors contain high-frequency graph signals, whereas low eigenvalues indicate that the corresponding eigenvectors contain low-frequency graph signals. For simplicity, suppose an undirected graph has $N$ nodes and each node only has a scalar feature so that the graph signal can be represented as a vector $\mathbf{x}\in\mathbb{R}^{N}$. Based on the eigendecomposition of the normalized graph Laplacian matrix of an undirected graph, the GFT and inverse GFT (IGFT) can be formulated in Equations (11) and (12), respectively,
$\widetilde{\mathbf{x}}=\mathbf{U}^{\prime}\mathbf{x};$ (11)
$\mathbf{x}=\mathbf{U}\widetilde{\mathbf{x}},$ (12)
where $\widetilde{\mathbf{x}}$ is the graph signal projected on the spectral domain.
Compared to spatial methods, the spectral graph signal $\widetilde{\mathbf{x}}$ contains more global information. Through GFT, the graph signal $\mathbf{x}$ is aggregated to produce its projection $\widetilde{\mathbf{x}}$ in the spectral domain according to the overall graph structure instead of the neighborhood information only. However, it is challenging to perform such GFT and IGFT based on eigendecomposition on directed graphs, whereas the spatial methods can naturally extend to directed graphs based on their neighborhood information passing design. Fortunately, a unified GSP framework can be applied to both undirected and directed graphs based on the magnetic Laplacian proposed by Zhang et al. (2021).
4.2 Magnetic Laplacian
Although the spatial methods can be more easily adapted to directed graphs compared to the spectral methods, they also have some drawbacks. Relying on the neighborhood information passing mechanism, spatial methods place more emphasis on out-going edges rather than in-going edges, thus limiting their ability to process the information on directed graphs (Zhang et al., 2021). The magnet Laplacian can give equal importance to different directions with its special design.
Suppose a directed graph $\mathcal{G}$ has $N$ nodes. In order to compute the magnet Laplacian $\mathbf{L}_{m}$, the symmetric variant of adjacency matrix $\mathbf{A}^{s}$ and the corresponding degree matrix $\mathbf{D}^{s}$ should be firstly defined in Equation (13) and Equation (14), respectively,
$\mathbf{A}^{s}=\frac{1}{2}(\mathbf{A}+\mathbf{A}^{\top});$ (13)
$\mathbf{D}^{s}=\text{diag}(\sum_{j=1}^{N}\mathbf{A}^{s}_{1j},\ldots,\sum_{j=1}^{N}\mathbf{A}^{s}_{Nj}).$ (14)
Besides, to capture the directional information, the phase matrix $\bm{\Theta}^{(q)}$ is constructed as shown in Equation (15),
$\bm{\Theta}^{(q)}=2\pi q(\mathbf{A}-\mathbf{A}^{\top}),$ (15)
where $q$ is a non-negative hyperparameter that determines how directional information is perceived and processed. Use $i$ to represent the imaginary unit. Based on $\mathbf{A}^{s}$ and $\bm{\Theta}^{(q)}$, the complex Hermitian adjacency matrix $\mathbf{H}^{(q)}$ is defined in Equation (16),
$\mathbf{H}^{(q)}=\mathbf{A}^{s}\odot\text{exp}(i\bm{\Theta}^{(q)}),$ (16)
where $\odot$ represents the element-wise multiplication between two matrices and $\text{exp}(\cdot)$ is also the element-wise exponential transformation applied to each element of the input matrix.
A complex Hermitian matrix $\mathbf{H}$ is a square matrix such that it is equal to its conjugate transpose: $\mathbf{H}=\mathbf{H}^{\dagger}$, which means $\mathbf{H}_{ij}=\overline{\mathbf{H}_{ji}}$ and $\overline{\mathbf{H}_{ji}}$ is the complex conjugate of element $\mathbf{H}_{ij}$. Furthermore, the unnormalized and normalized magnet Laplacians can be formulated in Equations (17) and (18), respectively,
$\mathbf{L}_{m}^{(q)}=\mathbf{D}^{s}-\mathbf{H}^{(q)};$ (17)
$\widetilde{\mathbf{L}}_{m}^{(q)}=\mathbf{I}-({(\mathbf{D}^{s})}^{-\frac{1}{2}}\mathbf{A}^{s}{(\mathbf{D}^{s})}^{-\frac{1}{2}})\odot\text{exp}(i\bm{\Theta}^{(q)}).$ (18)
Different $q$ values represent different handlings of the directionality of the graph. For example, if $q=0$, $\mathbf{H}^{(0)}=\mathbf{A}^{s}$. The directionality of the graph is eliminated through symmetrizing the asymmetric adjacency matrix. If $q=0.25$, $\mathbf{L}_{m}^{(0.25)}=-(\mathbf{L}_{m}^{(0.25)})^{\top}$ when there is only one directed edge between any two connected nodes of $\mathcal{G}$.
Take the normalized magnet Laplacian matrix $\widetilde{\mathbf{L}}_{m}^{(q)}$ as an example. Its eigendecomposition can be formulated in Equation (19) below:
$\widetilde{\mathbf{L}}_{m}^{(q)}=\mathbf{U}_{m}\bm{\Lambda}_{m}\mathbf{U}_{m}^{\dagger},$ (19)
where columns in $\mathbf{U}_{m}$ are still the orthonormal basis of eigenvectors of $\widetilde{\mathbf{L}}_{m}^{(q)}$ and $\mathbf{U}_{m}^{\dagger}$ is the conjugate transpose of $\mathbf{U}_{m}$. It has been proved that $\mathbf{U}_{m}$ and $\mathbf{U}_{m}^{\dagger}$ are complex valued while eigenvalues in $\bm{\Lambda}_{m}$ are all real non-negative numbers (Zhang et al., 2021). This allows easy extensions of classical GSP techniques to directed graphs. The GFT and IGFT for graph signal $\mathbf{x}\in\mathbb{R}^{N}$ based on normalized magnet Laplacian are formulated in Equations (20) and (21), respectively,
$\widetilde{\mathbf{x}}=\mathbf{U}_{m}^{\dagger}\mathbf{x},$ (20)
$\mathbf{x}=\mathbf{U}_{m}\widetilde{\mathbf{x}}.$ (21)
The magnet Laplacian offers a powerful tool to extend the GSP techniques to directed graphs. By adjusting the value of $q$, the magnet Laplacian can suit both the undirected and direct graph structures and can process the directional information in the graphs as needed. When $q>0$, the eigenvalue decomposition of the magnet Laplacian leverages the graph directional information by extending the graph Fourier bases to the complex domain, while the eigenvalues remain non-negative real numbers as meaningful graph signal frequency measurements. This is extremely useful in multivariate RV forecasting problems in that the volatility spillover effect between different markets can have different scales and directions (Diebold and Yilmaz, 2012). This means that, compared to an undirected or unweighted graph, a directed and weighted graph is more suitable to capture the volatility spillover network consisting of various global stock markets. In this case, compared to the GNNHAR model relying on spatial GNNs which are biased in processing directional information, models applying GSP techniques based on the magnet Laplacian are more powerful in learning the volatility dynamics (Zhang et al., 2021). The details on the construction of the directed and weighted volatility spillover graph are presented in the next section.
4.3 Volatility Spillover Graph Formulation
There are various ways to formulate the volatility spillover graph for global stock market indices based on their historical RV observations. For example, the Pearson correlation method and the Graphical LASSO (GLASSO) method (Friedman et al., 2007) can produce a symmetric matrix to represent the connections between the RV time series of different stock market indices. However, these methods cannot capture the imbalanced impact between influential stock markets and impressionable stock markets. On the other hand, the Diebold and Yilmaz (DY) framework can capture volatility spillovers from and to specific markets (Diebold and Yilmaz, 2012).
Suppose a covariance stationary multivariate RV time series $\{\mathbf{v}_{t}\in\mathbb{R}^{N}\}_{t=1}^{T}$ consists of $N$ different stock market indices. A $p$-order vector autoregressive model (VAR($p$)) for these time series can be represented as:
$\mathbf{v}_{t}=\sum_{i=1}^{p}\phi_{i}\mathbf{v}_{t-i}+\bm{\epsilon}_{t},\;\;\bm{\epsilon}_{t}\sim(0,\bm{\Sigma}_{\bm{\epsilon}}),$ (22)
where each $\phi_{i}$ is a scalar coefficient. This model can be rewritten in a moving average (MA) form:
$\mathbf{v}_{t}=\sum_{i=1}^{\infty}\mathbf{B}_{i}\bm{\epsilon}_{t-i},$ (23)
where ${\mathbf{B}_{i}}$ defined recursively by $\mathbf{B}_{i}=\sum_{j=1}^{p}\phi_{j}\mathbf{B}_{i-j}$ with the assumption that $\mathbf{B}_{0}=\mathbf{I}_{N\times N}$ and $\mathbf{B}_{i}=0$ for $i<0$. By performing a variance decomposition of the $H$-step-ahead forecast error, each element in the resulting matrix $\bm{\theta}^{g}(H)$ given by:
$\bm{\theta}_{ij}^{g}(H)=\frac{\sigma_{jj}^{-1}\sum_{h=0}^{H-1}(\mathbf{e}_{i}^{\top}\mathbf{B}_{h}\bm{\Sigma}_{\bm{\epsilon}}\mathbf{e}_{j})^{2}}{\sum_{h=0}^{H-1}(\mathbf{e}_{i}^{\top}\mathbf{B}_{h}\bm{\Sigma}_{\bm{\epsilon}}\mathbf{B}_{h}^{\top}\mathbf{e}_{i})},$ (24)
where $H$ is the forecast horizon and $\mathbf{e}_{i}$ is a selection vector of dimension $N$ with $1$ in the $i$-th position and zeros elsewhere. To standardize this matrix such that each row sums to $1$, the elements in the normalized matrix $\widetilde{\bm{\theta}}^{g}(H)$ can be defined as:
$\widetilde{\bm{\theta}}_{ij}^{g}(H)=\frac{\bm{\theta}_{ij}^{g}(H)}{\sum_{j=1}^{N}\bm{\theta}_{ij}^{g}(H)}.$ (25)
Furthermore, the net pairwise volatility spillovers $\overline{\bm{\theta}}^{g}(H)$ can be further calculated based on $\widetilde{\bm{\theta}}^{g}(H)$:
$\overline{\bm{\theta}}_{ij}^{g}(H)=\begin{cases}\widetilde{\bm{\theta}}_{ij}^{g}(H)-\widetilde{\bm{\theta}}_{ji}^{g}(H),&\mbox{$\widetilde{\bm{\theta}}_{ij}^{g}(H)\geq\widetilde{\bm{\theta}}_{ji}^{g}(H)$};\\ 0,&\mbox{$\widetilde{\bm{\theta}}_{ij}^{g}(H)<\widetilde{\bm{\theta}}_{ji}^{g}(H)$}.\end{cases}$ (26)
For this study, the DY framework is preferred because it incorporates directionality. Here, the hyperparameters $p=22$ to align the HAR model and $H$ is set according to specific forecast tasks. Besides, the net pairwise volatility spillover effect matrix $\overline{\bm{\theta}}^{g}(H)$ is applied because it can effectively highlight the dominant channels of volatility transmission, which allows researchers to pinpoint the markets that exert the most significant influence over others.
4.4 Model Formulation
In this section, the proposed vanilla GSPHAR (v-GSPHAR for short) and GSPHAR models are introduced. We first present the simpler and more basic v-GSPHAR model to demonstrate how GSP techniques are integrated into the HAR framework for RV forecasting. After that, a more flexible and sophisticated GSPHAR model is developed. The general workflow of the GSPHAR model is described in Figure 2 below:
Figure 2: GSPHAR model architecture.
Now we present the detailed step-by-step development of the proposed framework. Suppose that the multivariate RV time series $\{\mathbf{v}_{t}\in\mathbb{R}^{N}\}_{t=1}^{T}$ consist of the RV time series of $N$ different stock market indices from time $t=1$ to time $t=T$. In the v-GSPHAR model, we eliminate the directional information in the volatility spillover graph by setting the hyperparameter $q=0$ in the magnet Laplacian (18). Thus, the Hermitian adjacency matrix $\mathbf{H}^{(q)}$ and the symmetric adjacency matrix $\mathbf{A}^{s}$ become identical and they are constructed based on the net pairwise volatility spillovers $\overline{\bm{\theta}}^{g}(H)$:
$\mathbf{A}^{s}=\frac{1}{2}(\mathbf{A}+\mathbf{A}^{\top})=\frac{1}{2}(\overline{\bm{\theta}}^{g}(H)+\overline{\bm{\theta}}^{g}(H)^{\top}).$ (27)
The normalized magnet Laplacian matrix for the v-GSPHAR model is formulated as:
$\widetilde{\mathbf{L}}_{m}^{(0)}=\mathbf{I}-{(\mathbf{D}^{s})}^{-\frac{1}{2}}\mathbf{A}^{s}{(\mathbf{D}^{s})}^{-\frac{1}{2}}.$ (28)
$\widetilde{\mathbf{L}}_{m}^{(0)}$ is symmetric and positive semidefinite. Hence, it can be decomposed into real matrices through eigendecomposition as in Equation (10) and the multivariate RV time series can be projected to the spectral domain through the GFT as shown below:
$[\widetilde{\mathbf{v}}_{t-1},\widetilde{\mathbf{v}}_{t-2},\ldots,\widetilde{\mathbf{v}}_{t-22}]=\mathbf{U}_{m}^{\top}[\mathbf{v}_{t-1},\mathbf{v}_{t-2},\ldots,\mathbf{v}_{t-22}]$ (29)
where $\widetilde{\mathbf{L}}_{m}^{(0)}=\mathbf{U}_{m}\bm{\Lambda}_{m}\mathbf{U}_{m}^{\top}$. Hence, a set of basis-dependent univariate HAR models can be constructed in the graph Fourier domain:
${\widetilde{\mathbf{v}}_{t}}=\bm{\alpha}+\bm{\beta}_{d}\widetilde{\mathbf{v}}_{t-1}+\bm{\beta}_{w}\widetilde{\mathbf{v}}_{t-5:t-1}+\bm{\beta}_{m}\widetilde{\mathbf{v}}_{t-22:t-1},$ (30)
where $\widetilde{\mathbf{v}}_{t-n:t-1}=\frac{1}{n}\sum_{j=t-n}^{t-1}\widetilde{\mathbf{v}}_{j}$ is the average spectral RV data of each graph Fourier basis from time $t-1$ to time $t-n$. In addition, $\bm{\alpha}$, $\bm{\beta}_{d}$, $\bm{\beta}_{w}$ and $\bm{\beta}_{m}$ are all diagonal square matrices containing the graph Fourier basis–specific parameters. The forecast spectral RV values are projected back to the spatial domain to derive the final forecast output $\widehat{\mathbf{v}}_{t}$ through IGFT:
$\widehat{\mathbf{v}}_{t}=\mathbf{U}_{m}{\widetilde{\mathbf{v}}_{t}}.$ (31)
In this way, the model can be trained and the parameters $\bm{\alpha}$, $\bm{\beta}_{d}$, $\bm{\beta}_{w}$ and $\bm{\beta}_{m}$ can be estimated through the loss measured by any loss function with respect to the true RV value $\mathbf{v}_{t}$ and the forecast RV value $\widehat{\mathbf{v}}_{t}$.
In summary, the v-GSPHAR model constructs and trains a set of univariate HAR models in the graph Fourier domain. Compared to the VHAR and HAR-KS models, the number of parameters in the v-GSPHAR model increases linearly with the addition of more stock market indices. Furthermore, the volatility spillover effect is captured through GFT and its inverse (IGFT). This enables the model to perceive a broader impact range compared to methods that rely on spatial neighborhood information passing. However, this v-GSPHAR still holds the linearity assumption in model construction and cannot effectively capture the directional information in the volatility spillover graph. To address these limitations, we propose an improved GSPHAR model below.
Essentially, the improved GSPHAR model can be viewed as a combination of two v-GSPHAR models with learnable average pooling weights, integrated through a shallow neural network. To learn from the directional volatility spillover effect, the hyperparameter $q$ of the magnet Laplacian should be nonzero. Since the net pairwise volatility spillovers $\overline{\bm{\theta}}^{g}(H)$ is applied here, $q=0.25$ is chosen for better clarity and interpretability as $\mathbf{L}_{m}^{(0.25)}=-(\mathbf{L}_{m}^{(0.25)})^{\top}$ when only one directed edge exists between any two connected nodes of the volatility spillover graph. Since $q=0.25$, the eigendecomposition of the magnet Laplacian $\widetilde{\mathbf{L}}_{m}^{(0.25)}$ returns two eigenvector matrices, $\mathbf{U}_{m}$ and $\mathbf{U}_{m}^{\dagger}$, in the complex space and one eigenvalue matrix, $\bm{\Lambda}_{m}$, containing non-negative real numbers in its diagonal as shown in Equation (19). Thus, the multivariate RV time series input $[\mathbf{v}_{t-1},\mathbf{v}_{t-2},\ldots,\mathbf{v}_{t-22}]\in\mathbb{R}^{N\times 22}$ can be projected to the spectral domain through the GFT:
$[\widetilde{\mathbf{v}}_{t-1},\widetilde{\mathbf{v}}_{t-2},\ldots,\widetilde{\mathbf{v}}_{t-22}]=\mathbf{U}_{m}^{\dagger}[\mathbf{v}_{t-1},\mathbf{v}_{t-2},\ldots,\mathbf{v}_{t-22}].$ (32)
Here, each projected value in $\{\widetilde{\mathbf{v}}_{t-p}\}_{p=1}^{22}$ is complex valued and can be written as: $\widetilde{\mathbf{v}}_{t-p}=\widetilde{\mathbf{v}}^{\text{r}}_{t-p}+i\widetilde{\mathbf{v}}^{\text{i}}_{t-p}$. The projected RV data merges the volatility information between different stock markets under the guidance of the volatility spillover graph. The GSPHAR model is designed to process the information in both the real and imaginary domains in a way similar to the HAR family of models.
The HAR family of models always applies an equal-weight average to construct the volatility pattern features based on the historical RV data. This assumption can be potentially improved. For example, under the simple design of the univariate HAR model in Equation (2), $\text{RV}_{t-2}$ and $\text{RV}_{t-5}$ always exert the same influence on the forecast $\text{RV}_{t}$. This equal weighting may not be able to sufficiently account for the dynamics of financial markets where the influence of past volatility on future volatility typically diminishes over time. This diminishing influence not only exists between short-term, mid-term and long-term, but also possibly occurs within different periods. To handle such potential insufficiency, a set of graph Fourier basis-specific convolution filters with learnable convex weights is applied to aggregate past information in the mid-term and long-term:
$\begin{split}\widetilde{\mathbf{v}}_{i,t-n:t-1}&\;=\sum_{j=t-n}^{t-1}w_{i,j}\widetilde{\mathbf{v}}_{i,j},\\ s.t.&\;\sum_{j=t-n}^{t-1}w_{i,j}=1,w_{i,j}\geq 0\end{split}$ (33)
where $i\in{1,\ldots,N}$, each $w_{i,j}$ is a learnable non-negative weight in the convolution filter. The sum of all $w_{i,j}$ for each $i$ is equal to $1$ so that the weights are convex. In this way, the relative importance of each past RV observation can be easily depicted and compared with the weights in the original HAR model. The convolution filters are applied to the projected RV data so that it can aggregate (smooth) the spectral signals for each channel (time lag unit $j$) on each graph Fourier basis (each column in $\mathbf{U}_{m}$). Two distinct sets of convolution filters are applied to the mid-term and long-term spectral RV data, whereas the same set of convolution filters is used for both the real and imaginary parts of the spectral RV data within each time period. To analyze the overall impact of each time lag unit $j$, weights in each set of filters are averaged across all $N$ graph Fourier bases.
Hereafter, two distinct univariate HAR models are built for the real and imaginary domains, respectively. For clarity, let $\widetilde{\mathbf{V}}_{t-22:t-1}=[\mathbf{1},\widetilde{\mathbf{v}}_{t-1},\widetilde{\mathbf{v}}_{t-5:t-1},\widetilde{\mathbf{v}}_{t-22:t-1}]\in\mathbb{R}^{N\times 4}$ to denote a collection of intercept terms and past short-term, mid-term and long-term volatility patterns. The matrix $\mathbf{W}\in\mathbb{R}^{4\times 1}$ contains the linear transformation parameters used in the HAR model. Here, superscript ‘r’ is used to denote the real component and superscript ‘i’ represents the imaginary component. The same set of HAR coefficients is applied to different graph Fourier bases because the different lagged impact is believed to be well captured in the convolution filters defined in Equation (33). For each graph Fourier basis, its RV forecast in the real and imaginary domains can be formulated as:
$\begin{split}\widetilde{\mathbf{v}}^{\text{r}}_{t}&\;=\widetilde{\mathbf{V}}_{t-22:t-1}^{\text{r}}\mathbf{W}^{\text{r}},\\ \widetilde{\mathbf{v}}^{\text{i}}_{t}&\;=\widetilde{\mathbf{V}}_{t-22:t-1}^{\text{i}}\mathbf{W}^{\text{i}}.\end{split}$ (34)
Set $\widetilde{\mathbf{v}}_{t}=\widetilde{\mathbf{v}}^{\text{r}}_{t}+i\widetilde{\mathbf{v}}^{\text{i}}_{t}$, which is still a complex number. $\widetilde{\mathbf{v}}_{t}$ is projected back to the spatial domain through the IGFT:
$\overline{\mathbf{v}}_{t}=\mathbf{U}_{m}\widetilde{\mathbf{v}}_{t}.$ (35)
The complex valued $\overline{\mathbf{v}}_{t}$ is divided into its real and imaginary parts again: $\overline{\mathbf{v}}_{t}=\overline{\mathbf{v}}^{\text{r}}_{t}+i\overline{\mathbf{v}}^{\text{i}}_{t}$. The information of the real part and the imaginary part of $\overline{\mathbf{v}}_{t}$ is aggregated by a $3$ layer neural network $NN(\cdot)$ to output the final forecast: $\widehat{\mathbf{v}}_{t}=NN(\overline{\mathbf{v}}^{\text{r}}_{t},\overline{\mathbf{v}}^{\text{i}}_{t})$. The shallow neural network can merge the volatility information in the real and imaginary domains in a nonlinear way, which effectively enhances the expressiveness of the proposed GSPHAR model. Compared to the v-GSPHAR model, the improved GSPHAR model is more flexible in capturing the directional spillover effect and forecasting future volatility patterns at the cost of increasing parameters. Despite the complex model design, the growth rate of the parameters still remains linear with respect to the number of stock market indices included. This ensures that the GSPHAR model is not only powerful, but also scalable.
4.5 Dynamic Modeling
As mentioned in the Literature Review section, the pattern of interconnectedness among stock markets can change over time. However, there is a paucity of work on incorporating the dynamic volatility spillover effect into RV forecasting models. The proposed GSPHAR model offers potential solutions to handle the changing correlations in the global stock market. The GSPHAR model merges the volatility spillover effect into the RV data through GFT, in which the volatility network is precomputed and independent of the model training process. As long as the volatility network varies across different inputs, the model can become sensitive to the changes in volatility dynamics and adapt to the latest volatility patterns. However, how to change the volatility network with respect to the input RV data is an open question.
To address this, we propose a simple yet effective dynamic volatility spillover graph construction method based on both the DY framework and the Pearson correlation coefficient matrix of the input RV data. The DY framework is used to calculate the volatility network $\mathbf{A}^{\text{DY}}$ based on the whole in-sample dataset. $\mathbf{A}^{\text{DY}}$ can be interpreted as the average or overall volatility interconnection patterns during the in-sample period. The absolute Pearson correlation coefficient matrix $|\mathbf{A}^{\text{P}}|$ of the selected indices is calculated based on each RV input $\{\mathbf{v}_{t-p}\}_{p=1}^{22}$ of the HAR model. Specifically, two absolute Pearson correlation coefficient matrices, $|\mathbf{A}^{\text{P}}_{t-5:t-1}|$ and $|\mathbf{A}^{\text{P}}_{t-22:t-1}|$, are constructed based on the mid-term and long-term lagged RV input, respectively. For simplicity, the non-negative correlation coefficient matrices $|\mathbf{A}^{\text{P}}_{t-5:t-1}|$ and $|\mathbf{A}^{\text{P}}_{t-22:t-1}|$ are applied to the DY framework adjacency matrix $\mathbf{A}^{\text{DY}}$ in a element-wise basis to produce the updated adjacency matrix $\mathbf{A}$:
$\mathbf{A}=\rho|\mathbf{A}^{\text{P}}_{t-5:t-1}|\odot\mathbf{A}^{\text{DY}}+(1-\rho)|\mathbf{A}^{\text{P}}_{t-22:t-1}|\odot\mathbf{A}^{\text{DY}},$ (36)
where $\rho$ is a non-negative hyperparameter to balance the influence of mid-term historical RVs and long-term historical RVs when calculating $\mathbf{A}$. The outcome matrix $\mathbf{A}$ is leveraged to calculate the magnet Laplacian matrix, which is used in the GSPHAR model. The correlation coefficient matrices $|\mathbf{A}^{\text{P}}_{t-5:t-1}|$ and $|\mathbf{A}^{\text{P}}_{t-22:t-1}|$ can be regarded as a modification or adjustment applied to the magnitude of the overall volatility interconnection pattern $\mathbf{A}^{\text{DY}}$ such that the overall volatility network is adjusted towards the more recent RV data. A GSPHAR model that incorporates such dynamic designs is referred to as the dynamic-GSPHAR (d-GSPHAR) model.
5 Experiments
This section empirically evaluates the performance of the proposed v-GSPHAR, GSPHAR and d-GSPHAR models, using multivariate square root RV time series from $24$ stock market indices. The experiments are designed to compare these proposed models against baseline models to evaluate the validity and effectiveness of the GSP-based RV forecasting methods and to uncover new empirical insights into RV dynamics. The usefulness of the convolution filters in the GSPHAR model is also investigated. Furthermore, the d-GSPHAR model is compared with the GSPHAR model to determine the effect of incorporating volatility dynamics into the proposed framework.
5.1 Model Evaluation Tools
Various evaluation criteria are applied to evaluate the effectiveness of the proposed models. The evaluation process is conducted on an index-wise basis, as different markets operate under different economic conditions and investor behaviors. Especially, the accuracy of out-of-sample RV forecasts is measured through the commonly used Mean Absolute Error (MAE, calculation detail to be shown in the following Section 5.3). In addition, the Model Confidence Set (MCS) test (Hansen et al., 2011) and the Diebold-Mariano (DM) test (Diebold and Mariano, 2002) are applied to compare the accuracy of out-of-sample forecasts between the proposed models and baseline models to assess whether the improvement is significant or not. Similar to the input settings of the HAR models, the out-of-sample forecasting windows are set as $H=1,5,22$, corresponding to the short-term (daily), mid-term (weekly) and long-term (monthly) forecasting.
The software and hardware configurations are:
  • Software:
    • Python 3.10.14,
    • NumPy 1.26.4,
    • PyTorch 1.13.1+cu116.
  • Hardware:
    • Operating system: Windows 11,
    • CPU: 13th Gen Intel(R) Core(TM) i9-13900HX,
    • GPU: NVIDIA GeForce RTX 4080 Laptop GPU.
5.2 Benchmark Datasets and Baseline models
The benchmark dataset consists of the square root RV data of $24$ stock market indices. Among them, some indices are commonly used indices in previous RV forecast research (e.g., Liang et al. (2020)). The daily RV data is produced by Equation (1) based on the $5$-minute high-frequency returns. Given that multivariate RV forecast models are trained on RV observations on common trading days of the included stock market indices, the selection of these $24$ indices is intentionally focused on markets with more overlapping trading schedules. This approach ensures that the RV time series can be more continuous with more synchronized trading days across different markets, thus enhancing the reliability and comparability of the multivariate RV forecasts produced by different models. Besides, the whole dataset is scaled up by multiplying $100$.
A descriptive analysis of the scaled square-rooted data for each stock market index is reported in Table 11 in Appendix A. The Augmented Dickey-Fuller (ADF) test (MacKinnon, 1994; Cheung and Lai, 1995) is conducted and results reject the existence of the unit root in each time series at the $5\%$ significance level. This indicates that the time series data are stationary. As illustrated in the Volatility Spillover Graph Formulation section, the DY framework is applied to measure the net pairwise volatility spillover effect between different stock market indices. The net volatility spillover graph of the $24$ stock market indices calculated following Equation (26) is presented below. The nodes represent different indices and the numbers on the edges measure the net pairwise volatility spillover effect. Larger values indicate stronger volatility spillover effects while smaller values reflect weaker volatility transmission.
Figure 3: Net volatility spillover graph
The baseline models include the HAR, VHAR, HAR-KS and GNNHAR presented in the Literature Review section. Although the exact construction of some baseline models is not available, they are formulated according to the corresponding research papers. The relevant dataset and the code of the proposed models can be found at https://github.com/MikeZChi/GSPHAR.git.
5.3 Experiment Results
The general index-wise performance of each model is measured by the MAE calculated as follows:
$\text{MAE}_{i}=\frac{1}{T_{\text{out}}}\sum_{t=1}^{T_{\text{out}}}|\widehat{\mathbf{v}}_{i,t}-\mathbf{v}_{i,t}|,$ (37)
where $\text{MAE}_{i}$ is the MAE of the $i_{\text{th}}$ stock market index and $T_{\text{out}}$ denotes the duration of the out-of-sample dataset. $\widehat{\mathbf{v}}_{i,t}$ refers to the square root RV forecasting value of the $i_{\text{th}}$ index at time $t$ and $\mathbf{v}_{i,t}$ is the true square root RV value of the $i_{\text{th}}$ index at time $t$. The index-wise MAE results produced by each model with different forecast window $H$ are presented in Tables 1, 2 and 3.
Table 1: MAE comparison of models across different stock market indices in the short-term forecasting task. Minimum MAE values for each index are highlighted in bold. The second smallest MAE values for each index are underlined.
$H=1$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
[…]
Table 2: MAE comparison of models across different stock market indices in the mid-term forecasting task. Minimum MAE values for each index are highlighted in bold. The second smallest MAE values for each index are underlined.
$H=5$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
[…]
Table 3: MAE comparison of models across different stock market indices in the long-term forecasting task. Minimum MAE values for each index are highlighted in bold. The second smallest MAE values for each index are underlined.
$H=22$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
[…]
According to the MAE results in Tables 1, 2 and 3, it is clear that the GSPHAR model has improved RV forecasting results over all three forecast horizons. The GSPHAR model achieves the minimum MAE scores for $23$, $22$ and $21$ stock market indices in the short-term, mid-term and long-term RV forecasting tasks, respectively. Even when it does not achieve the best performance, the GSPHAR model still produces the second smallest MAE scores. These imply the consistent superiority of the GSPHAR model in different forecasting horizons. Besides, compared to the linear HAR family of models, including the HAR model, the VHAR and the HAR-KS model, the v-GSPHAR model performs better (achieves a smaller MAE score more often) in all forecasting tasks. This indicates that even the basic GSP techniques are more effective than the residual learning method of the baseline HAR models in integrating the volatility spillover effect. However, the v-GSPHAR model is inferior to the GSPHAR model and the GNNHAR model, especially in mid-term and long-term RV forecasting tasks. This implies that the linear expressiveness of the v-GSPHAR model could be insufficient in capturing the potential nonlinear volatility patterns, especially for the mid-term and long-term forecasts.
Although the MAE score tables have shown the improved performance of the GSPHAR model, further backtests are carried out to validate whether these improvements are statistically significant. At a specified significance level, the MCS test evaluates all tested models to determine a subset of models that has statistically superior forecasting accuracy (based on MAE) for each market index. On the 5% significance level, the results of the MCS test for the models under different forecasting horizons are presented in Tables 4, 5, and 6.
Table 4: The MCS test results table of the short-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at $5\%$ level.
$H=1$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
FCHI $\mathbf{0.066}$ 00 $120.0$ $\mathbf{1}$ $\mathbf{0.957}$
AEX 00 00 00 00 $\mathbf{0.522}$ $\mathbf{1}$
BFX $040.0$ $\mathbf{0.526}$ 00 $\mathbf{0.615}$ $\mathbf{1}$
STOXX50E $150.0$ $\mathbf{0.175}$ 00 $\mathbf{0.195}$ $\mathbf{1}$
IBEX $030.0$ 00 00 $\mathbf{0.263}$ $\mathbf{1}$
GDAXI $010.0$ 00 00 $0.001$ $\mathbf{1}$
AORD $0.019$ $\mathbf{0.108}$ $\mathbf{0.781}$ $0.004$ $\mathbf{0.561}$ $\mathbf{1}$
FTSE $0.002$ $\mathbf{0.144}$ $\mathbf{0.262}$ $0.001$ $\mathbf{0.837}$ $\mathbf{1}$
MXX $\mathbf{0.559}$ $0.015$ $\mathbf{0.155}$ $\mathbf{0.113}$ $\mathbf{0.559}$ $\mathbf{1}$
IXIC $\mathbf{0.056}$ 00 $0.009$ $0.007$ $0.015$ $\mathbf{1}$
SSMI $150.0$ 00 00 $\mathbf{0.111}$ $\mathbf{1}$
SPX $040.0$ 00 $0.001$ $0.004$ $\mathbf{1}$
RUT $0.040$ $\mathbf{0.761}$ $\mathbf{0.948}$ $0.012$ $\mathbf{0.827}$ $\mathbf{1}$
DJI $010.0$ $0.001$ $0.001$ $0.020$ $\mathbf{1}$
BSESN 00 00 00 00 $0.002$ $\mathbf{1}$
NSEI 00 00 00 00 $0.005$ $\mathbf{1}$
KS11 00 00 00 00 00 $\mathbf{1}$
BVSP $010.0$ 00 00 $\mathbf{0.096}$ $\mathbf{1}$
HSI $0.010$ $0.001$ $0.010$ $0.010$ $\mathbf{0.077}$ $\mathbf{1}$
KSE 00 00 00 00 00 $\mathbf{1}$
N225 00 $010.0$ 00 $0.001$ $\mathbf{1}$
SSEC 00 00 00 00 00 $\mathbf{1}$
OSEAX 00 00 $\mathbf{0.380}$ 00 00 $\mathbf{1}$
GSPTSE $160.0$ 00 $0.016$ $0.035$ $\mathbf{1}$
Table 5: The MCS test results table of the mid-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at $5\%$ level.
$H=5$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
FCHI 00 00 00 00 00 $\mathbf{1}$
AEX $040.0$ $010.0$ $0.004$ $\mathbf{1}$
BFX $120.0$ $0.012$ $0.001$ $0.012$ $\mathbf{1}$
STOXX50E 00 00 00 00 00 $\mathbf{1}$
IBEX 00 00 00 00 00 $\mathbf{1}$
GDAXI 00 00 00 00 00 $\mathbf{1}$
AORD $\mathbf{0.173}$ $0.009$ $\mathbf{1}$ $\mathbf{0.382}$ $\mathbf{0.284}$ $\mathbf{0.713}$
FTSE $280.0$ $\mathbf{0.097}$ $\mathbf{0.329}$ $\mathbf{0.05}$ $\mathbf{1}$
MXX $\mathbf{0.484}$ 00 $\mathbf{0.484}$ 00 $\mathbf{0.484}$ $\mathbf{1}$
IXIC $0.044$ $0.001$ $\mathbf{0.094}$ $\mathbf{0.277}$ $\mathbf{0.094}$ $\mathbf{1}$
SSMI $050.0$ $0.001$ $0.002$ $0.005$ $\mathbf{1}$
SPX $250.0$ $0.014$ $\mathbf{0.795}$ $0.008$ $\mathbf{1}$
RUT $\mathbf{0.187}$ $0.011$ $\mathbf{0.703}$ $\mathbf{0.566}$ $\mathbf{0.765}$ $\mathbf{1}$
DJI $250.0$ $0.046$ $\mathbf{0.587}$ $0.046$ $\mathbf{1}$
BSESN 00 00 00 00 00 $\mathbf{1}$
NSEI 00 00 00 00 00 $\mathbf{1}$
KS11 $0.004$ $0.005$ $040.0$ $0.005$ $\mathbf{1}$
BVSP $400.0$ $\mathbf{0.059}$ $\mathbf{1}$ $\mathbf{0.105}$ $\mathbf{0.767}$
HSI $190.0$ $190.0$ $\mathbf{0.056}$ $\mathbf{1}$
KSE $0.001$ $0.006$ $090.0$ $0.002$ $\mathbf{1}$
N225 $010.0$ $0.001$ $\mathbf{0.555}$ $0.001$ $\mathbf{1}$
SSEC 00 00 00 00 00 $\mathbf{1}$
OSEAX 00 00 $0.001$ $\mathbf{0.606}$ $0.004$ $\mathbf{1}$
GSPTSE $080.0$ $0.001$ $0.008$ $0.008$ $\mathbf{1}$
Table 6: The MCS test results table of the long-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at $5\%$ level.
$H=22$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
FCHI $090.0$ 00 $\mathbf{0.051}$ 00 $\mathbf{1}$
AEX $040.0$ 00 $0.021$ $0.001$ $\mathbf{1}$
BFX $050.0$ 00 $0.005$ $\mathbf{0.473}$ $\mathbf{1}$
STOXX50E $010.0$ 00 $0.007$ $0.001$ $\mathbf{1}$
IBEX $070.0$ 00 $070.0$ $\mathbf{1}$
GDAXI 00 00 00 $080.0$ $\mathbf{1}$
AORD $0.044$ $0.044$ $0.044$ $0.044$ $0.044$ $\mathbf{1}$
FTSE $070.0$ $0.001$ $0.013$ $0.013$ $\mathbf{1}$
MXX $\mathbf{0.594}$ $0.001$ $\mathbf{0.509}$ 00 $\mathbf{0.130}$ $\mathbf{1}$
IXIC $0.050$ $0.020$ $0.050$ $0.050$ $0.050$ $\mathbf{1}$
SSMI $030.0$ 00 $\mathbf{1}$ $0.003$ $\mathbf{0.887}$
SPX $100.0$ $0.002$ $\mathbf{0.200}$ $0.015$ $\mathbf{1}$
RUT $\mathbf{0.089}$ $0.005$ $0.036$ $0.013$ $\mathbf{0.167}$ $\mathbf{1}$
DJI $040.0$ 00 $\mathbf{0.128}$ $0.015$ $\mathbf{1}$
BSESN 00 00 00 $\mathbf{0.389}$ 00 $\mathbf{1}$
NSEI 00 00 00 $\mathbf{0.39}$ 00 $\mathbf{1}$
KS11 $040.0$ $0.008$ $0.021$ $0.004$ $\mathbf{1}$
BVSP $\mathbf{0.111}$ 00 $\mathbf{0.055}$ $\mathbf{0.111}$ $\mathbf{0.111}$ $\mathbf{1}$
HSI $\mathbf{0.271}$ 00 $0.002$ $\mathbf{0.106}$ $\mathbf{0.271}$ $\mathbf{1}$
KSE $190.0$ $\mathbf{0.201}$ $\mathbf{0.218}$ $0.035$ $\mathbf{1}$
N225 00 $030.0$ $\mathbf{0.175}$ $0.024$ $\mathbf{1}$
SSEC 00 00 00 $\mathbf{1}$ 00 $\mathbf{0.570}$
OSEAX $040.0$ 00 $0.008$ $0.008$ $\mathbf{1}$
GSPTSE $030.0$ 00 $\mathbf{1}$ 00 $\mathbf{0.920}$
According to the MCS test results, the GSPHAR model consistently remains in the MCS for all stock market indices in all short-term, mid-term and long-term forecasting tasks. On the other hand, other models stay in the MCS much less frequently.
The MAE results and the MCS test demonstrate that the GSPHAR model improved the RV forecasting performance compared to the other tested models. To further evaluate the RV forecasting results of the GSPHAR model across different forecast horizons, the one-sided DM test is applied to perform pairwise comparisons against the competing models.
Suppose the two candidate models are model 00 and model $1$. Their absolute forecast error series for the $i_{\text{th}}$ stock market index can be represented as $\{e^{(0)}_{i,t}\}_{t=1}^{T}$ and $\{e^{(1)}_{i,t}\}_{t=1}^{T}$. The distance between these two series of absolute forecast errors is measured as $d_{t}=e_{0,t}-e_{1,t}$. The test statistic $D$ of the DM test is calculated as follows:
$D=V_{d}^{-\frac{1}{2}}\overline{d},$ (38)
where $V_{d}$ is the autocorrelation-adjusted variance estimate of $\{d_{t}\}_{t=1}^{T}$ and $\overline{d}$ is the mean distance between the two absolute forecast error series: $\overline{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t}$. According to Equation (38), given that $V_{d}$ is non-negative, $D$ and $\overline{d}$ are both positive, negative or equal to 00. A positive $D$ indicates that, on average, $\{e_{1,t}\}_{t=1}^{T}$ is less than $\{e^{(0)}_{i,t}\}_{t=1}^{T}$ and model $1$ has higher forecast accuracy than model 00. The null hypothesis of the one-sided DM test is that model 00 and model $1$ are indifferent or model $1$ is inferior to model 00 in terms of forecast accuracy. When $D$ is sufficiently large, the DM test can reject the null hypothesis and conclude that model $1$ provides a significantly more accurate forecast than model 00. The results of the DM test are presented in Tables 7, 8 and 9. Here, for each DM test model $1$ represents the GSPHAR model, and model 00 represents other models.
Table 7: The DM test results table of the short-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at $5\%$ level.
$H=1$
HAR
vs GSPHAR
VHAR
vs GSPHAR
HAR-KS-RV
vs GSPHAR
GNNHAR
vs GSPHAR
v-GSPHAR
vs GSPHAR
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{2.414}$ $\mathbf{0.008}$ $\mathbf{9.383}$ $\mathbf{0}$ $\mathbf{3.332}$ $\mathbf{0}$ $\mathbf{7.360}$ $\mathbf{0}$ $-0.078$ $0.531$
AEX $\mathbf{7.361}$ $\mathbf{0}$ $\mathbf{27.021}$ $\mathbf{0}$ $\mathbf{14.82}$ $\mathbf{0}$ $\mathbf{16.156}$ $\mathbf{0}$ $1.404$ $0.080$
BFX $\mathbf{9.260}$ $\mathbf{0}$ $\mathbf{14.242}$ $\mathbf{0}$ $\mathbf{3.035}$ $\mathbf{0.001}$ $\mathbf{12.484}$ $\mathbf{0}$ $1.153$ $0.124$
STOXX50E $\mathbf{8.186}$ $\mathbf{0}$ $\mathbf{12.208}$ $\mathbf{0}$ $\mathbf{2.777}$ $\mathbf{0.003}$ $\mathbf{20.129}$ $\mathbf{0}$ $\mathbf{4.745}$ $\mathbf{0}$
IBEX $\mathbf{8.216}$ $\mathbf{0}$ $\mathbf{15.957}$ $\mathbf{0}$ $\mathbf{6.455}$ $\mathbf{0}$ $\mathbf{14.270}$ $\mathbf{0}$ $\mathbf{3.547}$ $\mathbf{0}$
GDAXI $\mathbf{10.085}$ $\mathbf{0}$ $\mathbf{37.947}$ $\mathbf{0}$ $\mathbf{6.074}$ $\mathbf{0}$ $\mathbf{22.738}$ $\mathbf{0}$ $\mathbf{7.345}$ $\mathbf{0}$
AORD $\mathbf{4.613}$ $\mathbf{0}$ $\mathbf{3.610}$ $\mathbf{0}$ $0.770$ $0.221$ $\mathbf{8.864}$ $\mathbf{0}$ $\mathbf{2.223}$ $\mathbf{0.013}$
FTSE $\mathbf{11.468}$ $\mathbf{0}$ $\mathbf{2.789}$ $\mathbf{0.003}$ $\mathbf{2.898}$ $\mathbf{0.002}$ $\mathbf{21.631}$ $\mathbf{0}$ $0.481$ $0.315$
MXX $\mathbf{2.034}$ $\mathbf{0.021}$ $\mathbf{6.491}$ $\mathbf{0}$ $\mathbf{2.58}$ $\mathbf{0.005}$ $\mathbf{15.513}$ $\mathbf{0}$ $0.980$ $0.164$
IXIC $\mathbf{3.711}$ $\mathbf{0}$ $\mathbf{6.044}$ $\mathbf{0}$ $\mathbf{6.441}$ $\mathbf{0}$ $\mathbf{16.923}$ $\mathbf{0}$ $\mathbf{6.551}$ $\mathbf{0}$
SSMI $\mathbf{5.491}$ $\mathbf{0}$ $\mathbf{27.432}$ $\mathbf{0}$ $\mathbf{14.34}$ $\mathbf{0}$ $\mathbf{13.679}$ $\mathbf{0}$ $\mathbf{3.621}$ $\mathbf{0}$
SPX $\mathbf{6.003}$ $\mathbf{0}$ $\mathbf{9.312}$ $\mathbf{0}$ $\mathbf{7.732}$ $\mathbf{0}$ $\mathbf{23.259}$ $\mathbf{0}$ $\mathbf{4.375}$ $\mathbf{0}$
RUT $\mathbf{5.559}$ $\mathbf{0}$ $0.779$ $0.218$ $0.083$ $0.467$ $\mathbf{11.084}$ $\mathbf{0}$ $0.979$ $0.164$
DJI $\mathbf{8.924}$ $\mathbf{0}$ $\mathbf{7.925}$ $\mathbf{0}$ $\mathbf{12.742}$ $\mathbf{0}$ $\mathbf{22.485}$ $\mathbf{0}$ $\mathbf{2.825}$ $\mathbf{0.002}$
BSESN $\mathbf{17.734}$ $\mathbf{0}$ $\mathbf{9.389}$ $\mathbf{0}$ $\mathbf{9.085}$ $\mathbf{0}$ $\mathbf{24.455}$ $\mathbf{0}$ $\mathbf{7.845}$ $\mathbf{0}$
NSEI $\mathbf{19.561}$ $\mathbf{0}$ $\mathbf{11.122}$ $\mathbf{0}$ $\mathbf{12.249}$ $\mathbf{0}$ $\mathbf{27.296}$ $\mathbf{0}$ $\mathbf{7.910}$ $\mathbf{0}$
KS11 $\mathbf{7.128}$ $\mathbf{0}$ $\mathbf{12.125}$ $\mathbf{0}$ $\mathbf{7.097}$ $\mathbf{0}$ $\mathbf{19.955}$ $\mathbf{0}$ $\mathbf{10.284}$ $\mathbf{0}$
BVSP $\mathbf{10.374}$ $\mathbf{0}$ $\mathbf{8.513}$ $\mathbf{0}$ $\mathbf{10.490}$ $\mathbf{0}$ $\mathbf{11.808}$ $\mathbf{0}$ $\mathbf{2.104}$ $\mathbf{0.018}$
HSI $\mathbf{5.070}$ $\mathbf{0}$ $\mathbf{13.739}$ $\mathbf{0}$ $\mathbf{2.889}$ $\mathbf{0.002}$ $\mathbf{6.463}$ $\mathbf{0}$ $\mathbf{3.854}$ $\mathbf{0}$
KSE $\mathbf{12.620}$ $\mathbf{0}$ $\mathbf{6.263}$ $\mathbf{0}$ $\mathbf{10.527}$ $\mathbf{0}$ $\mathbf{15.704}$ $\mathbf{0}$ $\mathbf{11.588}$ $\mathbf{0}$
N225 $\mathbf{17.899}$ $\mathbf{0}$ $\mathbf{3.805}$ $\mathbf{0}$ $\mathbf{8.245}$ $\mathbf{0}$ $\mathbf{20.918}$ $\mathbf{0}$ $\mathbf{8.258}$ $\mathbf{0}$
SSEC $\mathbf{14.961}$ $\mathbf{0}$ $\mathbf{6.670}$ $\mathbf{0}$ $\mathbf{14.393}$ $\mathbf{0}$ $\mathbf{34.431}$ $\mathbf{0}$ $\mathbf{14.830}$ $\mathbf{0}$
OSEAX $\mathbf{20.460}$ $\mathbf{0}$ $\mathbf{18.208}$ $\mathbf{0}$ $\mathbf{4.175}$ $\mathbf{0}$ $\mathbf{29.922}$ $\mathbf{0}$ $\mathbf{10.381}$ $\mathbf{0}$
GSPTSE $\mathbf{3.644}$ $\mathbf{0}$ $\mathbf{27.000}$ $\mathbf{0}$ $\mathbf{7.414}$ $\mathbf{0}$ $\mathbf{23.140}$ $\mathbf{0}$ $\mathbf{5.222}$ $\mathbf{0}$
Table 8: The DM test results table of the mid-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at $5\%$ level.
$H=5$
HAR
vs GSPHAR
VHAR
vs GSPHAR
HAR-KS-RV
vs GSPHAR
GNNHAR
vs GSPHAR
v-GSPHAR
vs GSPHAR
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{3.776}$ $\mathbf{0}$ $\mathbf{9.002}$ $\mathbf{0}$ $\mathbf{4.033}$ $\mathbf{0}$ $\mathbf{6.758}$ $\mathbf{0}$ $\mathbf{3.836}$ $\mathbf{0}$
AEX $\mathbf{2.858}$ $\mathbf{0.002}$ $\mathbf{7.192}$ $\mathbf{0}$ $\mathbf{4.365}$ $\mathbf{0}$ $\mathbf{6.206}$ $\mathbf{0}$ $\mathbf{3.125}$ $\mathbf{0.001}$
BFX $\mathbf{2.549}$ $\mathbf{0.005}$ $\mathbf{6.445}$ $\mathbf{0}$ $\mathbf{2.885}$ $\mathbf{0.002}$ $\mathbf{4.488}$ $\mathbf{0}$ $\mathbf{2.67}$ $\mathbf{0.004}$
STOXX50E $\mathbf{3.957}$ $\mathbf{0}$ $\mathbf{9.735}$ $\mathbf{0}$ $\mathbf{5.205}$ $\mathbf{0}$ $\mathbf{5.447}$ $\mathbf{0}$ $\mathbf{3.859}$ $\mathbf{0}$
IBEX $\mathbf{4.437}$ $\mathbf{0}$ $\mathbf{11.208}$ $\mathbf{0}$ $\mathbf{7.212}$ $\mathbf{0}$ $\mathbf{6.434}$ $\mathbf{0}$ $\mathbf{4.530}$ $\mathbf{0}$
GDAXI $\mathbf{3.631}$ $\mathbf{0}$ $\mathbf{9.271}$ $\mathbf{0}$ $\mathbf{4.579}$ $\mathbf{0}$ $\mathbf{6.203}$ $\mathbf{0}$ $\mathbf{4.085}$ $\mathbf{0}$
AORD $\mathbf{2.188}$ $\mathbf{0.014}$ $\mathbf{3.778}$ $\mathbf{0}$ $-0.522$ $0.699$ $1.24$ $0.107$ $1.045$ $0.148$
FTSE $\mathbf{3.611}$ $\mathbf{0}$ $\mathbf{7.421}$ $\mathbf{0}$ $\mathbf{3.059}$ $\mathbf{0.001}$ $1.337$ $0.091$ $\mathbf{3.340}$ $\mathbf{0}$
MXX $1.381$ $0.084$ $\mathbf{5.841}$ $\mathbf{0}$ $1.248$ $0.106$ $\mathbf{11.218}$ $\mathbf{0}$ $\mathbf{1.741}$ $\mathbf{0.041}$
IXIC $\mathbf{2.969}$ $\mathbf{0.001}$ $\mathbf{4.505}$ $\mathbf{0}$ $\mathbf{2.844}$ $\mathbf{0.002}$ $1.138$ $0.128$ $\mathbf{2.386}$ $\mathbf{0.009}$
SSMI $\mathbf{2.607}$ $\mathbf{0.005}$ $\mathbf{7.065}$ $\mathbf{0}$ $\mathbf{4.448}$ $\mathbf{0}$ $\mathbf{4.937}$ $\mathbf{0}$ $\mathbf{2.824}$ $\mathbf{0.002}$
SPX $\mathbf{3.897}$ $\mathbf{0}$ $\mathbf{5.530}$ $\mathbf{0}$ $\mathbf{3.805}$ $\mathbf{0}$ $0.269$ $0.394$ $\mathbf{4.414}$ $\mathbf{0}$
RUT $\mathbf{1.900}$ $\mathbf{0.029}$ $\mathbf{3.065}$ $\mathbf{0.001}$ $0.611$ $0.271$ $1.440$ $0.075$ $0.305$ $0.380$
DJI $\mathbf{4.193}$ $\mathbf{0}$ $\mathbf{4.990}$ $\mathbf{0}$ $\mathbf{2.985}$ $\mathbf{0.001}$ $0.665$ $0.253$ $\mathbf{3.736}$ $\mathbf{0}$
BSESN $\mathbf{5.572}$ $\mathbf{0}$ $\mathbf{8.433}$ $\mathbf{0}$ $\mathbf{8.362}$ $\mathbf{0}$ $\mathbf{8.526}$ $\mathbf{0}$ $\mathbf{4.202}$ $\mathbf{0}$
NSEI $\mathbf{6.111}$ $\mathbf{0}$ $\mathbf{9.776}$ $\mathbf{0}$ $\mathbf{9.058}$ $\mathbf{0}$ $\mathbf{9.041}$ $\mathbf{0}$ $\mathbf{4.504}$ $\mathbf{0}$
KS11 $\mathbf{4.403}$ $\mathbf{0}$ $\mathbf{2.834}$ $\mathbf{0.002}$ $\mathbf{4.536}$ $\mathbf{0}$ $\mathbf{6.473}$ $\mathbf{0}$ $\mathbf{3.646}$ $\mathbf{0}$
BVSP $\mathbf{2.941}$ $\mathbf{0.002}$ $\mathbf{5.760}$ $\mathbf{0}$ $\mathbf{2.332}$ $\mathbf{0.010}$ $-0.584$ $0.720$ $\mathbf{2.266}$ $\mathbf{0.012}$
HSI $\mathbf{2.520}$ $\mathbf{0.006}$ $\mathbf{4.747}$ $\mathbf{0}$ $\mathbf{2.139}$ $\mathbf{0.016}$ $\mathbf{4.679}$ $\mathbf{0}$ $\mathbf{1.959}$ $\mathbf{0.025}$
KSE $\mathbf{5.841}$ $\mathbf{0}$ $\mathbf{2.885}$ $\mathbf{0.002}$ $\mathbf{5.342}$ $\mathbf{0}$ $\mathbf{6.234}$ $\mathbf{0}$ $\mathbf{4.063}$ $\mathbf{0}$
N225 $\mathbf{6.927}$ $\mathbf{0}$ $\mathbf{4.755}$ $\mathbf{0}$ $\mathbf{4.412}$ $\mathbf{0}$ $0.736$ $0.231$ $\mathbf{4.988}$ $\mathbf{0}$
SSEC $\mathbf{5.813}$ $\mathbf{0}$ $\mathbf{5.702}$ $\mathbf{0}$ $\mathbf{7.178}$ $\mathbf{0}$ $\mathbf{5.112}$ $\mathbf{0}$ $\mathbf{6.451}$ $\mathbf{0}$
OSEAX $\mathbf{6.456}$ $\mathbf{0}$ $\mathbf{7.916}$ $\mathbf{0}$ $\mathbf{4.659}$ $\mathbf{0}$ $0.728$ $0.233$ $\mathbf{4.569}$ $\mathbf{0}$
GSPTSE $\mathbf{3.541}$ $\mathbf{0}$ $\mathbf{6.612}$ $\mathbf{0}$ $\mathbf{5.534}$ $\mathbf{0}$ $\mathbf{3.195}$ $\mathbf{0.001}$ $\mathbf{3.488}$ $\mathbf{0}$
Table 9: The DM test results table of the long-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at $5\%$ level.
$H=22$
HAR
vs GSPHAR
VHAR
vs GSPHAR
HAR-KS-RV
vs GSPHAR
GNNHAR
vs GSPHAR
v-GSPHAR
vs GSPHAR
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{3.056}$ $\mathbf{0.001}$ $\mathbf{6.550}$ $\mathbf{0}$ $\mathbf{4.696}$ $\mathbf{0}$ $\mathbf{2.066}$ $\mathbf{0.019}$ $\mathbf{3.741}$ $\mathbf{0}$
AEX $\mathbf{3.243}$ $\mathbf{0.001}$ $\mathbf{4.406}$ $\mathbf{0}$ $\mathbf{4.004}$ $\mathbf{0}$ $\mathbf{2.302}$ $\mathbf{0.011}$ $\mathbf{3.787}$ $\mathbf{0}$
BFX $\mathbf{2.278}$ $\mathbf{0.011}$ $\mathbf{3.548}$ $\mathbf{0}$ $\mathbf{2.855}$ $\mathbf{0.002}$ $\mathbf{2.890}$ $\mathbf{0.002}$ $0.690$ $0.245$
STOXX50E $\mathbf{3.488}$ $\mathbf{0}$ $\mathbf{6.351}$ $\mathbf{0}$ $\mathbf{4.853}$ $\mathbf{0}$ $\mathbf{2.968}$ $\mathbf{0.002}$ $\mathbf{3.426}$ $\mathbf{0}$
IBEX $\mathbf{2.729}$ $\mathbf{0.003}$ $\mathbf{7.057}$ $\mathbf{0}$ $\mathbf{5.014}$ $\mathbf{0}$ $\mathbf{3.364}$ $\mathbf{0}$ $\mathbf{3.519}$ $\mathbf{0}$
GDAXI $\mathbf{3.603}$ $\mathbf{0}$ $\mathbf{5.551}$ $\mathbf{0}$ $\mathbf{4.983}$ $\mathbf{0}$ $\mathbf{2.795}$ $\mathbf{0.003}$ $\mathbf{3.988}$ $\mathbf{0}$
AORD $\mathbf{2.529}$ $\mathbf{0.006}$ $\mathbf{2.119}$ $\mathbf{0.017}$ $\mathbf{2.913}$ $\mathbf{0.002}$ $\mathbf{2.724}$ $\mathbf{0.003}$ $\mathbf{3.180}$ $\mathbf{0.001}$
FTSE $\mathbf{3.048}$ $\mathbf{0.001}$ $\mathbf{3.396}$ $\mathbf{0}$ $\mathbf{3.567}$ $\mathbf{0}$ $\mathbf{3.254}$ $\mathbf{0.001}$ $\mathbf{2.718}$ $\mathbf{0.003}$
MXX $0.610$ $0.271$ $\mathbf{3.835}$ $\mathbf{0}$ $1.117$ $0.132$ $\mathbf{7.176}$ $\mathbf{0}$ $1.403$ $0.080$
IXIC $\mathbf{2.009}$ $\mathbf{0.022}$ $\mathbf{2.575}$ $\mathbf{0.005}$ $\mathbf{2.313}$ $\mathbf{0.010}$ $\mathbf{2.496}$ $\mathbf{0.006}$ $\mathbf{1.951}$ $\mathbf{0.026}$
SSMI $\mathbf{3.234}$ $\mathbf{0.001}$ $\mathbf{4.484}$ $\mathbf{0}$ $\mathbf{3.569}$ $\mathbf{0}$ $-0.149$ $0.559$ $\mathbf{3.476}$ $\mathbf{0}$
SPX $\mathbf{3.327}$ $\mathbf{0}$ $\mathbf{3.965}$ $\mathbf{0}$ $\mathbf{3.998}$ $\mathbf{0}$ $1.370$ $0.085$ $\mathbf{3.262}$ $\mathbf{0.001}$
RUT $\mathbf{1.667}$ $\mathbf{0.048}$ $\mathbf{3.054}$ $\mathbf{0.001}$ $\mathbf{2.246}$ $\mathbf{0.012}$ $\mathbf{2.920}$ $\mathbf{0.002}$ $1.387$ $0.083$
DJI $\mathbf{3.438}$ $\mathbf{0}$ $\mathbf{3.570}$ $\mathbf{0}$ $\mathbf{4.055}$ $\mathbf{0}$ $1.586$ $0.056$ $\mathbf{3.026}$ $\mathbf{0.001}$
BSESN $\mathbf{4.805}$ $\mathbf{0}$ $\mathbf{6.721}$ $\mathbf{0}$ $\mathbf{4.738}$ $\mathbf{0}$ $0.762$ $0.223$ $\mathbf{4.901}$ $\mathbf{0}$
NSEI $\mathbf{4.770}$ $\mathbf{0}$ $\mathbf{6.809}$ $\mathbf{0}$ $\mathbf{4.808}$ $\mathbf{0}$ $0.865$ $0.193$ $\mathbf{4.885}$ $\mathbf{0}$
KS11 $\mathbf{3.348}$ $\mathbf{0}$ $\mathbf{3.532}$ $\mathbf{0}$ $\mathbf{2.876}$ $\mathbf{0.002}$ $\mathbf{2.347}$ $\mathbf{0.009}$ $\mathbf{3.363}$ $\mathbf{0}$
BVSP $\mathbf{1.782}$ $\mathbf{0.037}$ $\mathbf{4.378}$ $\mathbf{0}$ $\mathbf{2.217}$ $\mathbf{0.013}$ $\mathbf{2.042}$ $\mathbf{0.021}$ $1.537$ $0.062$
HSI $1.461$ $0.072$ $\mathbf{3.727}$ $\mathbf{0}$ $\mathbf{2.915}$ $\mathbf{0.002}$ $\mathbf{2.840}$ $\mathbf{0.002}$ $1.300$ $0.097$
KSE $\mathbf{2.780}$ $\mathbf{0.003}$ $\mathbf{4.719}$ $\mathbf{0}$ $\mathbf{2.019}$ $\mathbf{0.022}$ $1.171$ $0.121$ $\mathbf{2.683}$ $\mathbf{0.004}$
N225 $\mathbf{4.172}$ $\mathbf{0}$ $\mathbf{3.109}$ $\mathbf{0.001}$ $\mathbf{4.467}$ $\mathbf{0}$ $1.416$ $0.078$ $\mathbf{3.036}$ $\mathbf{0.001}$
SSEC $\mathbf{3.797}$ $\mathbf{0}$ $\mathbf{4.616}$ $\mathbf{0}$ $\mathbf{5.387}$ $\mathbf{0}$ $-0.571$ $0.716$ $\mathbf{4.056}$ $\mathbf{0}$
OSEAX $\mathbf{3.656}$ $\mathbf{0}$ $\mathbf{4.902}$ $\mathbf{0}$ $\mathbf{3.825}$ $\mathbf{0}$ $\mathbf{2.520}$ $\mathbf{0.006}$ $\mathbf{3.274}$ $\mathbf{0.001}$
GSPTSE $\mathbf{4.156}$ $\mathbf{0}$ $\mathbf{4.627}$ $\mathbf{0}$ $\mathbf{5.507}$ $\mathbf{0}$ $-0.084$ $0.533$ $\mathbf{4.453}$ $\mathbf{0}$
According to the DM test results in Tables 7, 8 and 9, in all short-term, mid-term and long-term forecasting tasks, the GSPHAR model has significantly higher RV forecasting accuracy compared to all baseline models, including the HAR, VHAR, HAR-KS-RV and GNNHAR models, for the majority of, if not all, the stock market indices. Similar to the GNNHAR model, the GSPHAR model also applies both linear and nonlinear transformations in merging the asymmetric volatility information from both the real and imaginary domains. This indicates that the volatility spillover effect is directional and has a complex nonlinear mechanism. In addition, compared to the GNNHAR model, the application of the magnet Laplacian in the GSPHAR model allows more efficient and effective graph information propagation. The design of the GSPHAR model not only enables it to easily scale up to larger volatility networks, but also allows it to better capture the asymmetric relationship globally.
The GSPHAR model is also significantly better than the v-GSPHAR model for most stock market indices in different forecasting horizons. This indicates that using a positive $q$ in the magnet Laplacian and a shallow neural network to capture the directional effect through both the real and imaginary domains is effective, emphasizing the nonlinear, directional and asymmetric nature of the volatility spillover effect.
The HAR models utilize a simple average to aggregate past RV values, which implies equal importance of the lagged terms within the mid-term and long-term periods. However, the learnable weights of the convolutional filters of the GSPHAR model reveal that the importance of past RV information decays more dynamically rather than uniformly. According to Figure 4, the decaying pattern of the learned weights is different from the pattern of the prespecified weights in HAR models. The learned weights provide useful information on how the impact diminishes within the short-term, mid-term and long-term periods. Compared to the HAR model weights, the convolution filters put more weight on the $t-1$ RV data, the short-term volatility pattern. This implies that the autoregression in the volatility process is more reactive to recent market conditions. From time lag $2$ to time lag $5$, the learned weights of the GSPHAR model decrease from about $0.38$ to $0.1$ while the weights of HAR models remain constant at $0.2$. This suggests that the HAR model potentially underestimates the importance of the more recent volatility pattern at time lag $2$ and overestimates the importance of the volatility pattern from time lag $3$ to time lag $5$ in the mid-term period. In the long-term period, similar conclusions can be made. The learned weights firstly start above the constant weights from time lag $6$ to time lag $9$, then fluctuate around the constant weights from time lag $10$ to time lag $16$ and finally remain smaller than the constant weights from time lag $17$ to time lag $22$. Although the general pattern of the constant weights of the HAR models is reasonable in that they do put more emphasis on more recent volatility patterns, the stepwise decline between each time period and the stability within each time period overlook the nuanced contributions of different time lags. On the other hand, the learnable weights indicate a more sophisticated understanding of how the influence of past RV values fades with time in the volatility autoregression process and provide a more flexible representation that better captures the temporal relationships inherent in the data.
Figure 4: GSPHAR convolutional weights vs HAR model weights
Compared to the GNNHAR which uses a fixed optimizable hyperparameter $l$ to capture the local volatility interdependencies, the GSPHAR model only relies on the precomputed adjacency matrix to capture the volatility spillover effect. This adjacency matrix can be updated easily as its construction is independent of the model training process. The d-GSPHAR is proposed to incorporate the changing volatility spillover networks by updating the adjacency matrix with correlation coefficient matrices, as described in Equation (36). To evaluate the benefits of the d-GSPHAR in RV forecasting, the index-wise MAE results produced by GSPHAR and d-GSPHAR with different forecast horizons $H$ are presented in Table 10. Although the MAE results between GSPHAR and d-GSPHAR are close in the short-term forecasting task, the d-GSPHAR still delivers better performances in the mid-term and long-term RV forecasting.
Table 10: MAE comparison between GSPHAR and d-GSPHAR across different stock market indices and different forecasting horizons. Smaller MAE values for each index and each forecasting horizon are highlighted in bold.
Short-term ($H=1$)
Mid-term ($H=5$)
Long-term ($H=22$)
Index GSPHAR d-GSPHAR GSPHAR d-GSPHAR GSPHAR d-GSPHAR
FCHI $0.187$ $\mathbf{0.185}$ $0.236$ $\mathbf{0.234}$ $\mathbf{0.271}$ $0.274$
AEX $0.174$ $\mathbf{0.174}$ $\mathbf{0.222}$ $0.224$ $0.262$ $\mathbf{0.260}$
BFX $\mathbf{0.168}$ $0.170$ $0.211$ $\mathbf{0.211}$ $0.242$ $\mathbf{0.241}$
STOXX50E $0.221$ $\mathbf{0.217}$ $0.272$ $\mathbf{0.269}$ $0.308$ $\mathbf{0.308}$
IBEX $0.194$ $\mathbf{0.193}$ $0.240$ $\mathbf{0.239}$ $0.270$ $\mathbf{0.270}$
GDAXI $\mathbf{0.168}$ $0.172$ $0.222$ $\mathbf{0.217}$ $0.248$ $\mathbf{0.247}$
AORD $\mathbf{0.168}$ $0.169$ $0.204$ $\mathbf{0.203}$ $0.236$ $\mathbf{0.236}$
FTSE $0.221$ $\mathbf{0.221}$ $0.261$ $\mathbf{0.260}$ $0.303$ $\mathbf{0.301}$
MXX $0.168$ $\mathbf{0.167}$ $0.19$ $\mathbf{0.189}$ $0.211$ $\mathbf{0.208}$
IXIC $0.209$ $\mathbf{0.208}$ $0.282$ $\mathbf{0.279}$ $0.337$ $\mathbf{0.336}$
SSMI $0.128$ $\mathbf{0.126}$ $0.165$ $\mathbf{0.164}$ $\mathbf{0.210}$ $0.211$
SPX $\mathbf{0.188}$ $0.191$ $0.265$ $\mathbf{0.260}$ $0.319$ $\mathbf{0.318}$
RUT $0.197$ $\mathbf{0.197}$ $0.259$ $\mathbf{0.254}$ $0.300$ $\mathbf{0.299}$
DJI $\mathbf{0.183}$ $0.184$ $0.25$ $\mathbf{0.245}$ $0.305$ $\mathbf{0.304}$
BSESN $\mathbf{0.159}$ $0.161$ $0.199$ $\mathbf{0.198}$ $0.239$ $\mathbf{0.238}$
NSEI $\mathbf{0.159}$ $0.162$ $\mathbf{0.197}$ $0.198$ $0.242$ $\mathbf{0.241}$
KS11 $\mathbf{0.131}$ $0.134$ $0.167$ $\mathbf{0.165}$ $\mathbf{0.195}$ $0.196$
BVSP $\mathbf{0.197}$ $0.200$ $0.253$ $\mathbf{0.250}$ $0.319$ $\mathbf{0.318}$
HSI $\mathbf{0.147}$ $0.150$ $0.177$ $\mathbf{0.176}$ $0.204$ $\mathbf{0.200}$
KSE $0.187$ $\mathbf{0.187}$ $0.220$ $\mathbf{0.218}$ $0.255$ $\mathbf{0.255}$
N225 $\mathbf{0.173}$ $0.177$ $0.223$ $\mathbf{0.220}$ $0.248$ $\mathbf{0.247}$
SSEC $0.163$ $\mathbf{0.163}$ $0.199$ $\mathbf{0.199}$ $0.243$ $\mathbf{0.242}$
OSEAX $\mathbf{0.248}$ $0.253$ $0.289$ $\mathbf{0.287}$ $0.336$ $\mathbf{0.336}$
GSPTSE $0.124$ $\mathbf{0.124}$ $0.167$ $\mathbf{0.164}$ $0.206$ $\mathbf{0.206}$
6 Conclusion
In this paper, three novel multivariate volatility forecasting models, the v-GSPHAR, the GSPHAR and the d-GSPHAR models, are proposed. Key contributions are summarized below:
  • The proposed models are the first attempts to bring the GSP techniques to classical HAR models and achieve significantly better results compared to existing HAR-type baseline models. The proposed models do not handle the volatility spillover effect as a secondary adjustment for overlooking the interconnection between different assets. Rather, the proposed models incorporate the volatility spillover effect as an integral component of the model’s design, which ensures that this phenomenon is embedded to effectively capture the complex interdependencies within financial systems. Based on GSP techniques with the magnet Laplacian, these models extend the traditional HAR framework to capture volatility spillover effects across global financial markets and improve volatility forecasting accuracy. The empirical results demonstrate that the GSPHAR model consistently outperforms baseline models, including the univariate HAR model and its multivariate extensions, and the v-GSPHAR model in short-term, mid-term and long-term multivariate RV forecasting problems.
  • The proposed model provides new insights into the volatility dynamics. The convolution filters disclose the more detailed volatility decaying process. The decaying volatility autoregression does not only exist between the short-term, mid-term and long-term periods, but it also occurs within each historical period. Besides, the d-GSPHAR model novelly integrates the changing volatility spillover effect into the multivariate RV forecasting problem. The dynamic volatility network design allows the proposed model to capture the changing volatility spillover effect as detected in Diebold and Yilmaz (2012); Antonakakis et al. (2018); Gong et al. (2021), which also enhances the forecasting accuracy.
Compared to the existing multivariate HAR model extensions, the GSPHAR model offers a scalable and efficient solution for incorporating the volatility spillover effect without succumbing to common issues, such as insufficient scalability, time-consuming hyperparameter tuning, over-localization and fixed correlations. These findings suggest that the appropriate application of GSP techniques holds great potential for multivariate volatility forecasting and can provide valuable insights for market participants in financial risk management and decision-making. Furthermore, as mentioned in previous sections, how to efficiently and effectively integrate the dynamic volatility spillover effect into the GSPHAR-related forecasting models is an interesting problem for future research.
7 Disclosure of Interest
No conflict of interest is to be declared.
8 Disclosure of Funding
No funding was received.
9 Data Availability Statement
Data were downloaded from Oxford-man Institute’s realized library. The authors confirm that the data supporting the findings of this study are available within the supplementary materials of the paper.
Appendix A RV Time Series Data Statistics
Table 11: Descriptive Statistics and ADF Test Results for the RV Dataset of Various Stock Market Indices
Index Area Mean Std Dev Skewness Kurtosis ADF Stats ADF p-value
FCHI France 0.959 0.541 2.349 9.241 -5.430 0
AEX Netherlands 0.880 0.529 2.686 12.486 -7.681 0
BFX Belgium 0.816 0.446 2.604 12.192 -4.782 0
STOXX50E Eurozone 1.033 0.611 2.646 12.626 -6.699 0
IBEX Spain 1.025 0.557 2.775 16.836 -5.268 0
GDAXI Germany 1.001 0.598 2.461 9.298 -4.921 0
AORD Australia 0.604 0.379 4.218 35.781 -6.052 0
FTSE UK 0.884 0.548 3.479 25.224 -7.579 0
MXX Mexico 0.758 0.409 3.547 29.078 -7.802 0
IXIC USA 0.824 0.477 2.562 12.042 -8.098 0
SSMI Switzerland 0.766 0.471 4.253 33.192 -6.750 0
SPX USA 0.799 0.546 2.858 14.293 -8.293 0
RUT USA 0.762 0.451 2.953 15.755 -8.718 0
DJI USA 0.808 0.547 3.201 19.406 -8.508 0
BSESN India 0.939 0.551 4.814 54.345 -8.831 0
NSEI India 0.945 0.592 5.258 70.300 -8.937 0
KS11 South Korea 0.815 0.417 2.299 10.896 -4.590 0
BVSP Brazil 1.063 0.495 2.834 15.739 -8.743 0
HSI China (Hong Kong) 0.816 0.383 3.214 20.980 -4.657 0
KSE Pakistan 0.832 0.513 2.542 10.339 -5.933 0
N225 Japan 0.824 0.432 3.053 20.018 -9.530 0
SSEC China 1.085 0.682 2.445 9.384 -4.352 0
OSEAX Norway 0.945 0.588 5.532 82.454 -9.911 0
GSPTSE Canada 0.669 0.459 3.518 21.735 -5.183 0